# Generative or Discriminative? Revisiting Text Classification in the Era of Transformers

### v1

@@ 7 Limitations
While we conducted a thorough examination of generative and discriminative classifiers under standard i.i.d. assumptions, our findings may not generalize to scenarios involving distribution shifts, such as co-variate shift (Bickel et al., 2009) or concept shift (Roychowdhury et al., 2024). Our analysis was limited to traditional fine-tuning approaches, excluding emerging paradigms such as few-shot prompt-based in-context learning (Sun et al., 2023; Gupta et al., 2023) and parameter-efficient techniques like LoRA (Hu et al., 2022), which may uncover newer insights. Furthermore, our study focused exclusively on pure text classification, leaving the exploration of multi-modal scenarios involving tabular data (Pattisapu et al., 2025), images Lu et al. (2019), audio Kushwaha and Fuentes (2023), and other modalities for future work. An additional promising direction would be to investigate the use of pre-trained models, disentangling the effects of pre-training and fine-tuning (see Section 1), and to assess their effectiveness in cross-lingual transfer tasks. We leave these avenues for future exploration.
@@ Appendix A More Background and Related works
Calibration & Ordinality. Model calibration is crucial in classification, as it reflects how well predicted probabilities align with actual frequencies. Proper Scoring Rules (PSR) Merkle and Steyvers (2013) offer a theoretical basis for producing calibrated predictions: a scoring rule (i.e. loss function) is proper if its expected value is minimized only when predicted probabilities match the true distribution. All our modeling approaches—Generative (AR, MLM, Discrete Diffusion) and Discriminative (Encoder)—optimize proper scoring rules. GPT and MLM maximize likelihood, Discrete Diffusion optimizes a variational bound, and cross-entropy minimizes the KL-divergence between predicted and true distributions. Recent work Blasiok et al. (2023) shows that models trained with PSRs are often naturally calibrated when achieving low training loss, without requiring post-hoc calibration. This motivates us to empirically assess calibration across our models, as their differing architectures and objectives may still lead to varying calibration behaviors.

### v2

@@ 7 Limitations
While we conducted a thorough examination of generative and discriminative classifiers under standard i.i.d. assumptions, our findings may not generalize to scenarios involving distribution shifts, such as co-variate shift (Bickel et al., 2009) or concept shift (Roychowdhury et al., 2024). Our analysis was limited to traditional fine-tuning approaches, excluding emerging paradigms such as few-shot prompt-based in-context learning (Sun et al., 2023; Gupta et al., 2023) and parameter-efficient techniques like LoRA (Hu et al., 2022), which may uncover newer insights. Furthermore, our study focused exclusively on pure text classification, leaving the exploration of multi-modal scenarios involving tabular data (Pattisapu et al., 2025), images Lu et al. (2019), audio Kushwaha and Fuentes (2023), and other modalities for future work.
@@ Appendix A More Background and Related works
Calibration & Ordinality. Model calibration is crucial in classification, as it reflects how well predicted probabilities align with actual frequencies. Proper Scoring Rules (PSR) Merkle and Steyvers (2013) offer a theoretical basis for producing calibrated predictions: a scoring rule (i.e. loss function) is proper if its expected value is minimized only when predicted probabilities match the true distribution. All our modeling approaches—Generative (AR, MLM, Discrete Diffusion) and Discriminative (Encoder)—optimize proper scoring rules, but only ENC demonstrates consistently calibrated results because it optimizes a loss directly aligned with the classification task. Although the other paradigms also optimize strictly proper scoring rules that guarantee the lowest expected score when predictions match the target distribution, they optimize different objectives that do not perfectly align with classification. This mismatch in optimization targets explains the differences observed in calibration performance. GPT and MLM maximize likelihood, Discrete Diffusion optimizes a variational bound, and cross-entropy minimizes the KL-divergence between predicted and true distributions. Recent work Blasiok et al. (2023) shows that models trained with PSRs are often naturally calibrated when achieving low training loss, without requiring post-hoc calibration. This motivates us to empirically assess calibration across our models, as their differing architectures and objectives may still lead to varying calibration behaviors.
@@ Appendix B Implementation Details
Our analysis of inference latency reveals significant differences across architectures - refer to Table 5. While ENC and MLM demonstrate comparable inference speeds (requiring single forward passes), AR requires |K| forward passes for prediction, though this can be parallelized at the cost of increased computation. DIFF exhibits substantially higher latency, taking approximately 20-100x longer than ENC/MLM due to its iterative denoising process. Specifically, for a batch of 1024 examples (sequence length 128) on an A100 GPU, ENC and MLM take  0.03s for small models (3.3M params) to  1.3s for large models (120.4M params), while DIFF requires 16-25s across model sizes.
@@
Model Size Parameters ENC MLM AR DIFF
Small 3.3M $cellcolorlightgreen!500.027 \cellcolorlightyellow$!500.058 16.2
Medium 30.3M $cellcolorlightgreen!500.292 \cellcolorlightyellow$!500.510 20.52
Large 120.4M $cellcolorlightgreen!501.260 \cellcolorlightyellow$!502.070 24.8
Table 5: Model Size v/s Inference Latency (avg wall-clock time per batch in seconds)
@@ Appendix E More Main Results
Appendix F Results on Pretrained models
Figure 10: [Best viewed in color] Comparison of weighted-F1 scores between pretrained AR and ENC models ($\uparrow$ is better). (X-axis: sample size, Y-axis: weighted-F1 score)
Appendix G Revisiting the Bias-Variance Trade-off in Modern Text Classifiers
Figure 11: [Best viewed in color] Comparison of weighted F1-scores between training strategies ENC, DIFF, AR, $AR_{\text{pseudo}}$, and MLM across different sample sizes as model size increases. The plots highlight that the classical bias-variance trade-off phenomenon is only evident in small-data settings (i.e., lower sample sizes).
Figure 12: [Best viewed in color] Comparison of F1-Mean scores across different sample sizes and strategies as model size increases. The figure highlights that, in some cases, increasing training data can adversely affect performance.
The classical bias-variance trade-off, which predicts a U-shaped performance curve, has long been a foundational principle in machine learning (Geman et al., 1992). However, the emergence of the “double descent” (Nakkiran et al., 2019) phenomenon has challenged this view, demonstrating that test error can decrease as model complexity increases into the highly overparameterized regime (Belkin et al., 2019; Nakkiran et al., 2019).
In this work, we also conduct a fine-grained analysis of this behavior specifically in modern text classification, exploring how different model architectures AR, MLM, DIFF, and ENC with varying data and model sizes. The results depicted in Figure 11 illustrate that the classical bias-variance trade-off is predominantly observed in small-data settings (i.e., lower sample sizes), reinforcing the foundational principle in these regimes. Our findings offer critical empirical insights that nuance the prevailing theory. While we did not scale model size sufficiently to observe the full double descent curve, our results are consistent with the overparameterization regime. Specifically, as shown in Figure 11, the classical “inverted U” phenomenon is only evident in small-data settings. In contrast, for full-data settings, model performance is consistently non-decreasing with increasing model size, suggesting that larger models do not necessarily perform poorly when trained on ample data. This trend is particularly pronounced for the AR architecture, which exhibits a robustly non-decreasing performance curve across our experiments.
Moreover, Figure 12 visualizes the interplay between sample size and strategy more explicitly. We further investigate the “more data can hurt” phenomenon (Nakkiran et al., 2019), finding that while more data generally improves performance, a performance drop is occasionally observed in specific datasets such as sst5 and multiclasssentiment. This finding, illustrated in Figure 12, suggests that although the phenomenon exists, no universal dataset-specific pattern dictates its occurrence.
Appendix H Acknowledgments
The research and writing of this paper were led by Siva Rajesh Kasa, who oversaw the entire process from conception and experimental design through to implementation strategies, core experiments, analytical insights, and final manuscript approval. Karan Gupta managed end-to-end diffusion model experiments, owned the rebuttal and camera-ready cycles, and played a significant role in dataset details, implementation, preparing bias-variance analysis, and guiding the integration of calibration and ordinality sections. Sumegh Roychowdhury took the lead on noise robustness studies, compiled and presented the main results, generated key tables and figures, and coordinated the overall paper structure to ensure clarity and narrative cohesion across all sections. Ashutosh Kumar was instrumental in designing and implementing both the MLM and robustness experiments, contributed to pre-trained and encoder model evaluations, and co-authored substantial parts of the introduction, methodology, and related work sections, as well as leading dataset-centric analyses and diffusion studies. Yaswanth Biruduraju was responsible for all aspects of the calibration and ordinality experiments, including their conception, implementation, and analysis, ensuring these metrics were rigorously evaluated and reported. Santhosh Kumar Kasa built the core experimental infrastructure, particularly the discrete diffusion model implementation, ran experiments for both generative and discriminative paradigms, and maintained comprehensive codebase documentation to guarantee reproducibility. Nikhil Priyatam Pattisapu contributed the autoregressive and pseudo-generative methodologies, led the associated empirical studies on model and sample size scaling, authored relevant sections, and played a constructive role in team consensus around generative versus discriminative distinctions and multiple manuscript revisions. Arindam Bhattacharya anchored the theoretical framing by shaping the generative-discriminative comparison paradigm and was essential in literature review and conceptual foundations. Shailendra Agarwal and Vijay Huddar supported discussions, and provided manuscript feedbacks